# Regularized Adjusted Plus-Minus Models for Evaluating and Scouting Football (Soccer) Players using Possession Sequences

### v1

@@ 3.3 Generalized lasso
where $D$ is a matrix that allows for individual specification of lasso-type penalties. The penalization structure in equation (6) is termed generalized lasso (cf. e.g. Tibshirani and Taylor, 2011) and includes the ranking lasso as well as the more well-known fused lasso (Tibshirani et al., 2004).
Applying the generalized lasso penalty in the player evaluation setting is not completely straightforward. The aim is to cluster similar player contributions together, however as the structure of the regression matrix is $X=(X_{N_{inv}},X_{N_{of}})$, it does not make sense to consider the simple ranking lasso penalty. Instead, a more natural approach is to consider direct involvements separated from indirect ones. We first examine the structure of the matrix $D_{N_{inv}}$ for the coefficients corresponding to one of the sub-matrices of $X$, namely $X_{N_{inv}}$. The goal is to employ a penalty such that similar strength estimates are shrunken together, i.e. to apply the ranking lasso to the estimates of these columns. Thus the matrix $D_{N_{inv}}$ can be expressed as a block matrix
@@
containing three blocks $A^{\top}=(A_{1}^{\top},\dots,A_{N-1}^{\top})$, $B^{\top}=(B_{1}^{\top},\dots,B_{N-1}^{\top})$ and $C^{\top}=(C_{1}^{\top},\dots,C_{N-1}^{\top})$. Each $A_{i}$ is a matrix of zeroes of dimension $(N_{inv}-i)\times(i-1)$, each $B_{i}$ is a matrix (actually a vector) of ones of dimension $(N_{inv}-i)\times 1$ and each $C_{i}$ the negative identity matrix of dimension $N_{inv}-i$. The same penalization structure can be applied to the coefficients of indirect involvement, i.e. these are also shrunken according to the ranking lasso penalty and thus $D_{N_{inv}}=D_{N_{of}}$. In order to obtain the full penalization matrix $D$, the two sub-matrices $D_{N_{inv}}$ and $D_{N_{of}}$ need to be concatenated suitably:
@@
Indeed one can think of different approaches for the penalization matrix $D$, for example, if $D_{N_{of}}$ is the identity matrix, this would correspond to a classical lasso penalty on the indirect involvement coefficients. For the application to player evaluation, it makes however sense to stick to a structure where the coefficients of direct involvements in possessions are separated from the ones of indirect.
4 Computational results and comparison of approaches
@@ 4.1 Implementation
4.2 Results
4.2.1 Comparing ratings
In this section, the resulting ratings from the four approaches are analyzed and compared. We focus on the 2017/18 season of the Spanish La Liga and start by presenting the individual results and their implications and end with a comparison of the different models. In total, the methodology is applied to 556 players, and for each of the penalized regression approaches the estimates for direct involvement, indirect involvement as well as the sum thereof are analyzed.
Table 1: Top 5 of each position group (forward, midfielder, defender, goalkeeper) with regards to the sum of direct (Inv) and indirect involvement (OF) under the rating derived from ridge regression.
Score Rank
Position Player Inv OF Sum Inv OF Sum
Forward L. Messi 0.04088 0.03344 0.07432 1 1 1
L. Suárez 0.03545 0.02547 0.06092 2 3 2
Cristiano Ronaldo 0.03228 0.01284 0.04511 3 44 3
Lucas Vázquez 0.01474 0.02482 0.03956 19 4 4
Iago Aspas 0.03169 0.00347 0.03516 5 157 7
Midfielder Paulinho 0.00942 0.02667 0.03609 40 2 5
I. Rakitić 0.01143 0.02391 0.03534 32 8 6
Pablo Fornals 0.01553 0.01568 0.03121 17 31 10
T. Kroos 0.01840 0.00647 0.02486 8 115 19
André Gomes 0.00218 0.02249 0.02467 136 10 20
Defender Nacho 0.00822 0.02456 0.03278 56 5 8
Piqué 0.00758 0.02445 0.03204 62 6 9
Marcelo 0.00707 0.02192 0.02899 66 11 13
Jordi Alba 0.00995 0.01794 0.02789 37 19 14
D. Godín 0.00289 0.02159 0.02448 122 14 21
Goalkeeper M. ter Stegen 0.00365 0.02344 0.0271 105 9 15
K. Navas 0.00607 0.01682 0.02289 73 25 24
J. Oblak -0.00055 0.02188 0.02133 305 12 32
Neto -0.00413 0.01647 0.01234 484 27 76
Iago Herrerín -0.00106 0.00858 0.00752 346 84 124
We first discuss the results from ridge regression. Table 1 shows the top 5 players for each of the four broad position groups discussed in Section 3 based on the sum of the coefficient for direct involvement (Inv) and indirect involvement (OF). While using the sum of the coefficients as a measure of player strength is to some degree arbitrary, the resulting table provides interesting insight into the behavior of the penalty. Considering the rank columns in Table 1 it becomes clear that the direct involvement coefficients favor offensive players such as forwards, while defenders and goalkeepers score better on the indirect involvements. This behavior is expected due to the way we set up our models. As the outcome variable measures whether a possession ends in a goal or not, clearly forwards should have a more direct impact on it since scoring goals is their main job. On the other hand, the main task of defenders and goalkeepers is to protect the team from receiving goals, which is rather measured through the indirect involvements. Thus, in order to measure the value of forwards and midfielders it may be more accurate to analyze their rank based on direct involvement as shown in Table 2. In the same manner, it may be more accurate to rank defenders and goalkeepers based on their indirect involvement coefficient as presented in Table 3.
Table 2: Top 10 forwards and midfielders with regards to direct involvement (Inv) in possessions under the rating derived from ridge regression.
Score Rank
Position Player Inv OF Sum Inv OF Sum
Forward L. Messi 0.04088 0.03344 0.07432 1 1 1
L. Suárez 0.03545 0.02547 0.06092 2 3 2
Cristiano Ronaldo 0.03228 0.01284 0.04511 3 44 3
Maxi Gómez 0.03223 -0.01875 0.01348 4 544 68
Iago Aspas 0.03169 0.00347 0.03516 5 157 7
A. Griezmann 0.02288 0.00327 0.02615 6 163 16
Munir 0.01868 -0.01013 0.00856 7 493 110
C. Bacca 0.01833 -0.00198 0.01636 9 350 50
Rodrigo 0.01799 0.00070 0.01868 11 237 40
G. Bale 0.01668 0.00025 0.01692 13 255 46
Midfielder T. Kroos 0.01840 0.00647 0.02486 8 115 19
D. Wass 0.01815 -0.00181 0.01635 10 343 51
Samu Castillejo 0.01686 0.00495 0.02182 12 133 29
P. Sisto 0.01635 0.00335 0.01971 15 161 37
Pablo Fornals 0.01553 0.01568 0.03121 17 31 10
P. Hernández 0.01457 -0.01252 0.00205 20 512 211
L. Modrić 0.01392 0.00031 0.01423 23 251 66
Portu 0.01364 0.00063 0.01426 24 239 65
I. Rakitić 0.01143 0.02391 0.03534 32 8 6
Mikel Oyarzabal 0.01076 0.00094 0.0117 35 230 81
Table 3: Top 10 defenders and goalkeeper with regards to indirect involvement (OF) in possessions under the rating derived from ridge regression.
Score Rank
Position Player Inv OF Sum Inv OF Sum
Defender Nacho 0.00822 0.02456 0.03278 56 5 8
Piqué 0.00758 0.02445 0.03204 62 6 9
J. Murillo -0.00177 0.02394 0.02217 389 7 25
Marcelo 0.00707 0.02192 0.02899 66 11 13
D. Godín 0.00289 0.02159 0.02448 122 14 21
Nélson Semedo 0.00325 0.01993 0.02318 113 15 23
J. Giménez 0.00217 0.01959 0.02176 138 16 30
Víctor Ruíz 0.00453 0.01950 0.02403 94 17 22
L. Hernández -0.00389 0.01855 0.01466 480 18 61
Jordi Alba 0.00995 0.01794 0.02789 37 19 14
Goalkeeper M. ter Stegen 0.00365 0.02344 0.0271 105 9 15
J. Oblak -0.00055 0.02188 0.02133 305 12 32
K. Navas 0.00607 0.01682 0.02289 73 25 24
Neto -0.00413 0.01647 0.01234 484 27 76
Fernando Pacheco -0.00645 0.01235 0.00590 530 52 143
Iago Herrerín -0.00106 0.00858 0.00752 346 84 124
Rubén Blanco -0.00247 0.00771 0.00523 432 95 155
Y. Bounou -0.00492 0.00653 0.00161 504 114 222
Kiko Casilla -0.00222 0.00567 0.00345 416 127 185
C. Pantilimon -0.00167 0.00467 0.00300 386 137 196
Next, the results from the group lasso regression are analyzed. For this exercise, the columns are split into 8 groups. The players’ direct and indirect involvements are grouped into the 4 player positions, resulting in 4 direct involvement groups and 4 indirect involvement groups. Table 4 is the analog of Table 1, i.e. it displays the top 5 players for each of the four positions under the group lasso penalty using the sum of the coefficients of direct (Inv) and indirect (OF) involvements.
Table 4: Top 5 of each position group (forward, midfielder, defender, goalkeeper) with regards to the sum of direct (Inv) and indirect involvement (OF) under the rating derived from group lasso regression.
Score Rank
Position Player Inv OF Sum Inv OF Sum
Forward Iago Aspas 0.74980 0.02308 0.77288 1 162 1
L. Messi 0.61590 0.15345 0.76934 3 9 2
Cristiano Ronaldo 0.68337 -0.01686 0.66650 2 389 3
L. Suárez 0.59205 0.02455 0.61660 4 155 4
Juanmi 0.49060 0.00323 0.49382 7 256 5
Midfielder Portu 0.40491 0.04977 0.45468 11 75 6
D. Wass 0.38755 -0.01856 0.36898 12 394 10
T. Inui 0.28829 0.03389 0.32218 18 119 15
Zurutuza 0.12923 0.13300 0.26223 65 11 23
Samu Castillejo 0.22036 0.0269 0.24725 30 141 27
Defender Nacho 0.05847 0.16819 0.22666 111 5 31
Víctor Ruíz 0.04315 0.16532 0.20847 131 7 35
Aritz Elustondo 0.11675 0.08441 0.20116 74 36 38
H. Moreno 0.08787 0.10561 0.19348 85 19 41
José Ángel 0.05719 0.10806 0.16525 113 17 50
Goalkeeper K. Navas 0.11321 0.0232 0.13642 76 159 72
J. Cillessen -0.01402 0.06396 0.04995 343 58 143
Toño 0.00000 0.04072 0.04072 226 103 157
Oier 0.00766 0.02601 0.03367 194 147 170
Jaume 0.01657 0.01000 0.02657 182 218 188
The general effect of the group lasso penalty is similar to the ridge regression, i.e. offensive players (forwards and midfielders) load higher on the coefficient of direct involvement and defensive players (defenders and goalkeepers) exhibit a tendency to load higher on the coefficient of indirect involvement. However, it seems that the effects are more pronounced, especially for the direct involvements. This is reasonable since the idea of the group lasso is to shrink whole groups and thus less relevant groups, such as the defensive players, are penalized more severely. Figure 1 analyses the relationship between the ridge estimates and the group lasso estimates in more detail. In terms of correlation between the two rating methods, it seems that there is the most agreement for the coefficients of direct involvements rather than the indirect involvements. Furthermore, the ratings of the offensive players are more highly correlated than the ones for the defensive players. Overall the coefficients seem to be similar in both types of involvement in possessions as in general, the correlation is greater than 0.75. The sum of the coefficients however exhibits the least correlation between the two ratings, suggesting that it may be worth studying different approaches of how to combine the two individual contributions to possession than simply summing them.
Figure 1: Correlation between the ratings from ridge regression and group lasso regression. Correlation of the scores (x-axis) and the ranks (y-axis) are displayed where each of the 4 position groups (forwards, midfielders, defenders, goalkeepers) are colored differently and each of the involvement types (Inv, OF, Sum) is displayed by a different symbol.
The results from the exclusive lasso are presented in Table 5 and Table 6. Again the same 8 groups as for the group lasso penalty are considered. As the exclusive lasso performs lasso-like shrinkage within the group but selects at least one relevant representative from each group, most of the players are shrunken to zero. However, the general pattern is similar to the two previous approaches. When analyzing direct involvement as in Table 5 offensive players score better than defensive ones. In fact, only one of the 3 selected defenders has a positive coefficient estimate and both estimates of the selected goalkeepers are negative. The coefficient estimates for the indirect involvement are displayed in Table 6. In this case, most of the defenders obtain positive values and the top player according to indirect involvement to possessions is a goalkeeper. Furthermore, the estimates of the selected midfielders are also positive, however, all of them are considered to primarily play in the center with strong defensive roles. Only two forwards are selected with only Lionel Messi obtaining a positive estimate of indirect involvement. Interestingly, Messi is also the only player selected with regards to both, involvement categories, and at the same time being the top selected player with regards to direct involvement. This suggests that his appearance on the field has a substantial impact on the game. As already mentioned, the results from the exclusive lasso are not optimal for rating or ranking players as only a few relevant players are not set exactly to zero. This precludes the possibility of comparing any set of players that may be of interest. In Appendix B, we discuss how the exclusive lasso penalty may be used for scouting players by defining groups of players that are of interest. In Section 4.3 we discuss a different approach for defining groups which is more useful for comparing the validity of the developed ratings.
Table 5: Selected players from each position group with regards to direct involvement (Inv) in possessions under the exclusive lasso penalty.
Score Rank
Position Player Inv OF Inv OF
Forward L. Messi $0.302$ $0.002$ $1$ $12$
L. Suárez $0.109$ $0.000$ $3$ $13$
Cristiano Ronaldo $0.043$ $0.000$ $6$ $13$
Maxi Gómez $0.033$ $0.000$ $7$ $13$
Midfielder T. Kroos $0.128$ $0.000$ $2$ $13$
Samu Castillejo $0.067$ $0.000$ $4$ $13$
D. Wass $0.064$ $0.000$ $5$ $13$
Defender Barragán $0.005$ $0.000$ $8$ $13$
Gabriel -$0.028$ $0.000$ $553$ $13$
R. Rosales -$0.112$ $0.000$ $556$ $13$
Goalkeeper Cuéllar -$0.048$ $0.000$ $554$ $13$
Arrizabalaga -$0.076$ $0.000$ $555$ $13$
Table 6: Selected players from each position group with regards to indirect involvement (OF) in possessions under the exclusive lasso penalty.
Score Rank
Position Player Inv OF Inv OF
Forward L. Messi $0.302$ $0.002$ $1$ $12$
J. Calleri $0.000$ -$0.409$ $9$ $556$
Midfielder Paulinho $0.000$ $0.172$ $9$ $2$
André Gomes $0.000$ $0.093$ $9$ $5$
I. Rakitić $0.000$ $0.031$ $9$ $8$
Tomás Pina $0.000$ $0.002$ $9$ $11$
Defender J. Murillo $0.000$ $0.149$ $9$ $3$
Nacho $0.000$ $0.128$ $9$ $4$
Víctor Ruíz $0.000$ $0.042$ $9$ $7$
Marcelo $0.000$ $0.023$ $9$ $9$
Luis Hernández $0.000$ -$0.008$ $9$ $552$
Goalkeeper J. Oblak $0.000$ $0.189$ $9$ $1$
Neto $0.000$ $0.050$ $9$ $6$
Fernando Pacheco $0.000$ $0.018$ $9$ $10$
Raúl Lizoáin $0.000$ -$0.009$ $9$ $553$
M. Dmitrović $0.000$ -$0.016$ $9$ $554$
Rubén $0.000$ -$0.020$ $9$ $555$
The last penalized regression approach discussed in this work is the generalized lasso. As described in Section 3.3, the shrinkage employed tries to cluster the coefficient contributions of players into groups of equal strength. The results for the penalty employed on the direct involvement coefficients are displayed in Figure 2. In order to be coherent, the players are again divided into the usual 4 position groups, although the implicit clustering of the generalized lasso does not account for position groups. The main message is consistent with the previous approaches. Offensive players perform better with regard to direct involvement than defensive players. However, most of the players are grouped into a specific sub-group of equal strength (total rank group 25 of Figure 2). Interestingly, the players selected to deviate from the average in terms of strength, i.e. the ones not grouped into the common strength group, differ from the ones selected under the exclusive lasso penalty, although not fundamentally. Furthermore, it can be noticed that the coefficients for all players are negative. This suggests that overall the contribution of average players to possessions is negative and thus also the contribution of strong players is shrunken towards a negative value. The results from the generalized lasso regularization employed on the coefficients of indirect involvement are represented in Figure 3. Once again a common strength group can be identified, which contains most of the players (total rank group 19 of Figure 3). In this case, mostly defenders and central respectively defending midfielders obtain higher scores, however, goalkeepers are primarily grouped into the cluster of average players. Finally, it has to be mentioned that the ranking is from its conceptual design to some extent similar to the exclusive lasso and thus also not ideal for ranking players in general. Nevertheless, both models provide interesting insights and the results suggest that strong (resp. weak) players can be isolated from the mass. However, the distinction between players that are average is difficult even though it may theoretically be possible within the ridge or group lasso framework.
@@ 4.2.1 Comparing ratings
Figure 3: Results for each of the 4 position groups with regards to indirect involvement (OF) under the generalized lasso penalty. Coefficient groups are represented by their coefficient estimate (x-axis) and their total rank (y-axis). Groups containing only one player are marked by the name of the players.
@@ 4.3 Validity of ratings
As the main goal of data-driven approaches for evaluating football players is to provide an objective view of the strengths of players, it is equally necessary to objectively analyze whether a novel metric is doing a good job. This task is not straightforward as there is no ground truth for the strength of players and the opinions of practitioners and experts may vary substantially on that matter. One naive idea would be to compare the results from the novel ratings with commonly used key performance indicators such as goal scored, assists, or even more statistical metrics such as expected goals (xG, c.f. e.g. Robberechts and Davis, 2020). However, the main reason for deriving more elaborated player rating systems is to observe player strengths beyond classical performance measures, and thus a more reliable approach is desired. This section is dedicated to providing a statistical evaluation of the rating methods described in this work.
@@
One criterion for a reasonable evaluation system of player strength discussed by Franks et al. (2016) and Hvattum and Gelade (2021) can be summarized under the term “relevance’ ’. A metric is relevant, if there is a relationship between the metric and an outcome of interest. In terms of football, the most interesting outcome is the result of a match or the number of matches won. Thus, a relevant metric should be able to establish a predictive relationship between the player strengths and the result of a match. That is, if the ratings developed under our framework are relevant, the prediction of match results should be improved by taking into account the strengths of the players involved in these matches.
@@
In order to test the relevancy of the ratings derived in this paper, we follow an approach that is conceptually similar to the validity checks in Hvattum and Gelade (2021). However, as we only work with data from the 2017/18 season of the Spanish La Liga the framework is adjusted similarly to the one in Bajons (2023). Match outcome data of said season is divided into a training set, containing the first 280 matches of the season, and a test set, containing the remaining 100 matches. For each match, a strength estimate of each of the competing teams is given as the average over the strength values of players playing for the team. As the penalized regression models contain two different measures of strength of each player (Inv and OF), we use two distinct methods for concatenating them. First, the sum of both effects is used as discussed in Section 4.2.1. Second, the estimates are standardized such that in both cases they range from 0 (indicating the weakest player) to 1 (indicating the strongest player). Then we simply take the average of them to determine a player’s strength. Furthermore, we adapt the grouping for the exclusive lasso. As the exclusive lasso by design attempts to set most of the players in a particular group to zero, considering only the four position groups as is done in Section 4.2.1 results in too many players set exactly to zero. A more interesting grouping for predicting match results is to consider the four position groups for each specific team that is playing. In this way, the algorithm selects the strongest (and/or possibly weakest) players on each of the position groups for each team.
@@
For prediction of match results two state-of-the-art models are considered. First, a bivariate Poisson (BP) model as discussed by Karlis and Ntzoufras (2003), which models the number of goals by the home and away team jointly based on covariates. Second, an ordered logistic regression (OLR) model as proposed in Arntzen and Hvattum (2021), which uses an ordered categorical variable with three levels as response, namely home win ($H$), draw ($D$) and away win. In particular, for both prediction models, only one covariate is used in order to analyze the respective outcomes, namely the difference in strengths of the home and away team. The strength of a team is hereby defined as the average value of the ratings as derived from each variant of our framework. The BP and OLR models are trained on the training set and the predictive performance is evaluated on the test set using proper scoring rules as suggested by Arntzen and Hvattum (2021) and Hvattum and Gelade (2021). First, the Brier score
@@
for each match in the test set is computed. Second, the informational loss is computed, which is given as
@@
where $p_{i}$ is the model’s probability of outcome $i=H,D,A$ and $d_{i}=1$, if match ended with outcome $i$ and 0 otherwise. That is, for all matches in the test set, the Brier score loss and the informational loss are obtained. Thus, the evaluation of a model can be done by considering the averages of these scores over the test period. Furthermore, differences in the resulting averages can be statistically tested by using a two-tailed paired $t$-test.
@@
Table 7 provides the results of the predictive performance when using the respective strength estimate as a covariate. Furthermore, three reference or baseline models are included. The model termed Baseline'' is a simple model without covariate and theELO’ ’ model is derived by using ELO team ratings as taken from http://clubelo.com/. The “PCV’ ’ model uses the $PCV$ metric described in Bajons (2023). Details for the calculation of the $PCV$ value as well as a comparison of the metric with the ratings described here are found in Appendix A. The best-performing model in both prediction frameworks and under both evaluation criteria seems to be the group lasso model, however, the ridge regression model is able to keep up, even once emerging on the top. The exclusive lasso follows these two models closely and the generalized lasso performs in general worse than the aforementioned models. This result is not unexpected as we already discussed that both these models are not necessarily optimal for deriving a rating as a lot of players are merged to the average and there is no possibility to distinguish them. However, all of these models seem to perform better in terms of validity than the two reference models. Interestingly, the PCV [^4] In the initial work, Bajons (2023) uses a slightly different data setup than the one used in this paper. More details are found in Appendix A. and the ELO model both perform only slightly better than the baseline model and are outperformed by the regularized regression models. Figure 4 displays a heatmap of the p-values of individual paired $t$-tests performed on the predictive losses of the 100 matches in the test set. The results from the $t$-test indicate that at least under the OLR model the best performing models of the ridge, group lasso, and exclusive lasso penalty significantly outperform the reference models (Baseline and ELO) at the 5 % level. For the bivariate Poisson model, the differences to the baseline model are only significant at the 10 % level, but the pattern is similar. These results suggest that the regularized regression models provide a rating that from an objective viewpoint captures the important parts of the game quite well.

### v2

@@ 3.3 Generalized lasso
where $D$ is a matrix that allows for individual specification of lasso-type penalties. The penalization structure in equation (6) is termed generalized lasso (cf.~e.g. Tibshirani and Taylor, 2011) and includes the ranking lasso as well as the more well-known fused lasso (Tibshirani et al., 2004).
Applying the generalized lasso penalty in the player evaluation setting is not completely straightforward. The aim is to cluster similar player contributions together, however as the structure of the regression matrix is $X=(X_{N_{inv}},X_{N_{of}})$, it does not make sense to consider the simple ranking lasso penalty. Instead, a more natural approach is to consider direct involvements separated from indirect ones. We first examine the structure of the matrix $D_{N_{inv}}$ for the coefficients corresponding to one of the sub-matrices of $X$, namely $X_{N_{inv}}$. The goal is to employ a penalty such that similar strength estimates are shrunken together, i.e.~to apply the ranking lasso to the estimates of these columns. Thus the matrix $D_{N_{inv}}$ can be expressed as a block matrix
@@
containing three blocks $A^{\top}=(A_{1}^{\top},\dots,A_{N-1}^{\top})$, $B^{\top}=(B_{1}^{\top},\dots,B_{N-1}^{\top})$ and $C^{\top}=(C_{1}^{\top},\dots,C_{N-1}^{\top})$. Each $A_{i}$ is a matrix of zeroes of dimension $(N_{inv}-i)\times(i-1)$, each $B_{i}$ is a matrix (actually a vector) of ones of dimension $(N_{inv}-i)\times 1$ and each $C_{i}$ the negative identity matrix of dimension $N_{inv}-i$. The same penalization structure can be applied to the coefficients of indirect involvement, i.e.~these are also shrunken according to the ranking lasso penalty and thus $D_{N_{inv}}=D_{N_{of}}$. In order to obtain the full penalization matrix $D$, the two sub-matrices $D_{N_{inv}}$ and $D_{N_{of}}$ need to be concatenated suitably:
@@
Indeed one can think of different approaches for the penalization matrix $D$, for example, if $D_{N_{of}}$ is the identity matrix, this would correspond to a classical lasso penalty on the indirect involvement coefficients. For the application to player evaluation, it makes however sense to stick to a structure where the coefficients of direct involvements in possessions are separated from the ones of indirect. In this way, this penalization approach allows for the selection of clusters of players with equal strength for on-ball and off-ball involvements separately. This may be of interest, when the goal is to identify players of similar strength and playing style compared to a reference player.
4 Computational details and results
@@ 4.1 Implementation
4.2 Comparing ratings
We provide a comparison of the penalization methods described in section 3. The aim here is to highlight idiosyncrasies as well as commonalities of each approach. A more thorough analysis containing specific rankings, player analyses, and further results can be found in the supplementary material of this work.
First, we compare the ratings obtained from ridge regression to the ones obtained from the group lasso. For the latter, the columns are split into eight groups. The players’ direct and indirect involvements are grouped into four player positions (forward, midfielder, defender, goalkeeper), resulting in four direct involvement groups and four indirect involvement groups. In general, the effect of the group lasso penalty is similar to the ridge regression as is visually confirmed by Figure 1. Offensive players (forwards and midfielders) load higher on the coefficient of direct involvement (Inv) and defensive players (defenders and goalkeepers) tend to load higher on the coefficient of indirect (on-field) involvement (OF) as is shown in the boxplots of Figure 1. However, it seems that the effects are more pronounced, especially for the direct involvements. This is reasonable since the idea of the group lasso is to shrink whole groups and thus less relevant groups, such as the defensive players, are penalized more severely. The left part of Figure 1 analyses the relationship between the ridge estimates and the group lasso estimates in more detail. In terms of correlation between the two rating methods, it seems that there is the most agreement for the coefficients of direct involvements rather than the indirect involvements. Furthermore, the ratings of the offensive players are more strongly correlated than the ones for the defensive players. Overall, the coefficients seem to be similar in both types of involvement in possessions with correlations of 0.7 or greater for all position groups. When trying to obtain a universal score for players, an intuitive approach would be to consider the sum of the coefficients from direct and indirect involvements. However, as seen in Figure 1 (left panel), the correlation between the two ratings is the lowest for the sum, suggesting that it may be worth studying different approaches of combining the two individual contributions to possession rather than simply summing them. Finally, interesting analyses can be done when comparing the Inv and OF for specific players and position groups, which we discuss in more detail in the supplement.
Figure 1: Comparison of the results from the ridge penalty and the group lasso penalty. Left: Correlation between the score of each group on the x-axis vs. correlation between the rank of each group on the y-axis. Groups are categorized by type of involvement (Inv, OF, sum) and position (FW, MF, DF, GK) Right: Boxplots of the estimates of each player in the 4 position groups under the group lasso and ridge penalty, separated by type of involvement.
Next, we discuss the results from the exclusive lasso. To allow for comparability, we use the same grouping as for the group lasso. Recall, that the exclusive lasso performs lasso-like shrinkage within each group but selects at least one relevant representative from each group. Figure 2 displays all selected players from the four direct involvement (Inv) as well as the four indirect involvement (OF) groups. The general pattern is similar to the two previous approaches. The offensive groups load stronger on Inv, with more players selected and higher coefficients (left panel of Figure 2). For OF, defensive players and groups score higher (right panel of Figure 2). As mentioned in Section 3.2, the exclusive lasso has its advantage in scouting, which would require a different setup of the groups. We refer the interested reader to the supplementary materials, where we present specific applications and discuss benefits and pitfalls when using this penalization variant.
Figure 2: Estimates of the players selected from the exclusive lasso penalty. Left: Players selected from the 4 direct involvement (Inv) position groups. Right: Players selected from the 4 indirect involvement (OF) position groups.
Finally, we analyze the results from the generalized lasso. In contrast to the group lasso and the exclusive lasso, this penalty allows for implicit grouping of players. Figure 3 shows the results for direct involvement coefficients. For comparability, we separate the players again into the usual four position groups. The main message is again consistent with the previous approaches: Offensive players perform better with regard to direct involvement than defensive players. We omit the results for indirect involvement and refer again to the supplementary material. Interestingly, the grouping formed is rather dense, i.e. outstanding players (positively and negatively) form their own groups while average players are bundled in one large group, suggesting that the distinction between these players is difficult even though theoretically possible within the ridge or group lasso framework.
@@ 4.3 Validity of ratings
Having derived a statistically motivated rating, an equally important task is to objectively analyze its validity. One criterion for a reasonable evaluation system of player strength discussed by Franks et al. (2016) and Hvattum and Gelade (2021) can be summarized under the term “relevance’ ’. A metric is relevant, if there is a relationship between the metric and an outcome of interest. In terms of football, the most interesting outcome is the result of a game. Thus, a relevant metric can establish a predictive relationship between a player’s strengths and the result of a match.
@@
The validity checks conducted in this paper follow the conceptual idea of Hvattum and Gelade (2021). %As we only work with data from the La Liga data from 2017/18, the framework is adjusted similarly to the one in Bajons (2023). In particular, match outcome data of the 2017/18 season is divided into a training set containing the first 280 matches and a test set containing the remaining 100. For each, a strength estimate of the competing teams is derived as the average over the strength values of players playing for the team. As the penalized regression models contain two different measures of player strength (Inv and OF), two distinct methods for concatenating them are applied. First, the sum of both effects is used as discussed in Section 4.2. Second, the estimates are standardized such that they range from 0 (indicating the weakest player) to 1 (indicating the strongest player) for the two contributions. Then, we average them to determine a player’s strength. Furthermore, we adapt the grouping for the exclusive lasso. As the exclusive lasso by design attempts to set most of the players in a particular group to zero, considering only the four position groups, as is done in Section~4.2, results in impractically strong sparsity. A more representative grouping for predicting match results considers the four position groups for all teams in our data. In this way, the algorithm selects the strongest (and/or possibly weakest) players on each position group for every team.
@@
Two state-of-the-art models are considered for the prediction of match results. First, a bivariate Poisson (BP) model as discussed by Karlis and Ntzoufras (2003), which models the number of goals by the home and away team jointly based on covariates. Second, an ordered logistic regression (OLR) model as proposed in Arntzen and Hvattum (2021), which uses an ordered categorical variable with three levels as response, namely home win ($H$), draw ($D$) and away win. In particular, for both prediction models, only one covariate is used to analyze the respective outcomes, namely the difference in strengths of the home and away teams. The strength of a team is defined as the average value of the ratings as derived from each variant of our framework. Predictive performance for the models is evaluated on the test set using proper scoring rules as suggested by Arntzen and Hvattum (2021) and Hvattum and Gelade (2021). We compute the Brier score
@@
and the informational loss
@@
where $p_{i}$ is the model’s probability of outcome $i=H,D,A$ and $d_{i}=1$, if match ended with outcome $i$ and 0 otherwise. That is, for all matches in the test set, the Brier score loss and the informational loss are obtained. The evaluation of a model is done by considering the averages of these scores over the test period. Furthermore, differences in the resulting averages can be statistically tested using a two-sided paired $t$-test.
@@
Table 1 provides the results of the predictive performance when using the respective strength estimate as a covariate. Furthermore, three reference models are included. The model termed Baseline'' is a simple model without covariate, and theELO’ ’ model is derived by using ELO team ratings as taken from http://clubelo.com/. The “PCV’ ’ model uses the $PCV$ metric described in Bajons (2023). Details for the calculation of the $PCV$ value as well as a comparison of the metric with the ratings described here are found in the supplementary material for this work. The best-performing model in both prediction frameworks and under both evaluation criteria seems to be the group lasso model. Although the ridge regression model is able to keep up, it seems that there is value in taking the positional information into account. The exclusive lasso follows these two models closely, and the generalized lasso performs in general worse than the aforementioned models. This result is not unexpected as both these models are not necessarily optimal for deriving a rating but rather have strengths in different use cases. However, all of these models perform better in terms of validity than the three reference models. Interestingly, the PCV and the ELO model both perform only slightly better than the baseline model and are outperformed by the regularized regression models. Figure 4 displays a heatmap of the p-values of individual paired $t$-tests performed on the predictive losses of the 100 matches in the test set. The results from the $t$-test indicate that at least under the OLR model the best performing models of the ridge, group lasso, and exclusive lasso penalty significantly outperform the reference models (Baseline and ELO) at the 5 % level. For the bivariate Poisson model, the differences to these models are only significant at the 10 % level, but the pattern is similar. These results suggest that the regularized regression models provide a rating that from an objective viewpoint captures the relevant parts of the game well.